A report in LTO stated that the average age of taxis in the Philippines is 9 years. An operations manager of a large taxi company selects a sample of 40 taxis and finds the average age of the taxis is 8.2 years. The o if the population is 2.3 years. At a = 0.05, can it be concluded that the average age of the taxis in his company is less than the national average?


1
Expert's answer
2022-03-31T07:42:55-0400

"\mu=9, \sigma=2.3,n=40,\bar{x}=8.2,\alpha=0.05."

Null and alternative hypotheses:

"H_0:\mu\geq9;\\
H_1:\mu<9."


Because "\sigma" is known and "n=40>30," we can use the z-test.

"z=\cfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\cfrac{8.2-9}{2.3/\sqrt{40}}=-2.20."

In z-table, the area corresponding to "z=-2.20" is 0.0139. Because the test is a left-tailed test, the P-value is equal to the area to the left of "z=-2.20," so, "P=0.0139."


Because the P-value is less than "\alpha" =0.05, we reject the null hypothesis, there is enough evidence at the 5% level of significance to support the claim that the average age of the taxis in this company is less than the national average.